OpenStudy (anonymous):

Use the half angle formula to find the exact value of cos(-3pi/8)

OpenStudy (anonymous):

\[\cos 2\theta = \cos^2 \theta - \sin^2 \theta = 1- 2\sin^2\theta = 2\cos^2 \theta - 1\] Try to use any one of the above formula and get the answer...

OpenStudy (anonymous):

Firstly you will use this formula : \[\cos (- \theta) = \cos \theta \] So, \[\cos(-3 \pi /8) = \cos (3 \pi /8)\] Now use the above identity...

OpenStudy (anonymous):

thats actually for the double angle formula isnt it? Half angle is +-sq.rt. 1-cos(a)/2?

OpenStudy (anonymous):

Yes buddy you are given cos2x Just find the cosx or sinx from the identities I have given...

OpenStudy (anonymous):

ok after plugging root 2/2 into the formula I got \[(\sqrt{2-\sqrt{2}})/4\]

OpenStudy (anonymous):

Buddy, thanks I appreciate you for doing this.. You are absolutely correct...